log[3 root x] = 4/3log[x-6]
Answer by stanbon(75887)

log[3 root x] = 4/3log[x-6]
----
log(3*sqrt(x)] = (4/3)log(x-6)
----
log(3sqrt(x)) = log(x-6)^(4/3)
------
3sqrt(2) = (x-6)^(4/3)
=====
Comment:I don't think this is the problem you intended.
==========
Cheers,
Stan H.